The fox population in a certain region has a relative growth rate of 9 percent per year. It is estimated that the population in the year 2000 was 29,900 foxes.
a) Find an equation that models the population t years after 2000 ( t=0 for 2000).
A =
b) Use the equation from part (a) to estimate the fox population in the year 2010. Round to the nearest fox.
= _____ foxes

Answer by Boreal(15235)

F=Fo(1.09)^t
F=29900(1.09)^t
where t=year-2000
F(2010)=29900(1.09)^10=70784 foxes